# Iterative Approach to Reconstructing Neural Disparity Fields from Light-field Data

### v1

@@ IV-A2 Specific scene experiments
To gain a comprehensive understanding of the performance of the proposed method, we set up isolated challenges in specific scenarios: backgammons, dots, pyramids, and stripes. The Backgammon scene evaluated the algorithm’s capability in handling fine structures and occlusion boundaries. The Dots scene investigated the impact of different levels of camera noise on the reconstruction of objects of various sizes. The Pyramids scene tested the algorithm’s performance on convex, concave, circular, and planar geometrical structures, and the Stripes scene specifically assessed the influence of texture and contrast on occlusion boundary recognition. We conducted comparative experiments using the proposed method, MV, LF_OCC, SPO, AMN, EPI_ORM[20], LLFNet[22], and FastLFnet[23]. The results are presented in Fig. 4.
@@
Figure 4: Visual comparison of disparity estimation methods across different scenes. From top to bottom, rows correspond to Backgammon, Dots, Pyramids, and Stripes scene. For each scene, the top row shows the estimated disparity, and the bottom row shows the corresponding MSE. Results (b)-(h) are derived from original data in [47].
@@
TABLE I: Qualitative comparisons of results on synthetic datasets, presented with BP01(BadPix0.01), BP03(BadPix0.03), BP07(BadPix0.07) and MSE ($\times$ 100) metrics.
MV LF_OCC SPO AMN EPI_ORM LLFNet FastLFNet Ours
Backgammon BP01 65.842 88.839 49.940 43.599 34.316 65.975 39.839 54.878
BP03 18.547 39.826 8.639 13.354 7.238 34.337 11.409 16.378
BP07 8.924 19.006 3.782 6.913 3.988 14.853 5.138 6.656
MSE 13.225 21.587 4.587 12.287 3.411 8.058 3.986 8.929
Dots BP01 83.519 65.762 58.079 21.658 65.707 67.417 68.146 4.538
BP03 59.654 23.552 35.068 18.428 47.932 41.971 41.109 4.485
BP07 45.587 5.822 16.274 18.428 36.098 20.134 21.169 4.380
MSE 7.262 3.301 5.238 2.936 14.480 6.159 3.407 5.433
Pyramids BP01 85.392 74.144 79.206 32.485 19.063 49.817 22.188 32.675
BP03 9.824 20.495 6.263 5.643 1.301 6.495 2.193 2.853
BP07 0.776 3.172 0.861 0.449 0.324 1.059 0.620 0.129
MSE 0.051 0.098 0.043 0.023 0.016 0.035 0.018 0.016
Stripes BP01 76.372 53.669 21.879 13.499 55.135 79.080 63.400 8.389
BP03 62.417 21.257 15.460 7.966 13.943 47.050 32.594 8.053
BP07 46.192 18.408 14.987 5.835 6.871 23.837 9.442 7.518
MSE 12.173 8.131 6.955 2.128 1.744 2.390 0.984 2.518
Average BP01 77.781 70.604 52.276 27.810 43.556 65.572 48.393 25.12
BP03 37.611 26.283 16.358 11.348 17.604 32.463 21.826 7.942
BP07 25.370 11.602 8.976 7.906 11.820 14.971 9.092 4.671
MSE 8.178 8.279 4.206 4.344 4.913 4.161 2.099 4.224
Paremeteres (M) 9.1 1.8 1.4 0.5
@@
Figures 4 (b)-(h) correspond to the different algorithms. From the experimental results, it is evident that the proposed method successfully reconstructs disparities in various challenging scenarios. In the Backgammon scene, the proposed method performed remarkably well in handling complex textures, tilted surfaces, and textureless smooth regions, particularly when compared with MV, LF_OCC, and SPO, although it failed to accurately estimate disparities at the upper-left crevice. In the Dots scene, the proposed method reconstructs more foreground points in the upper-left corner compared to the lower-right corner, indicating a certain robustness to noise; however, as the noise increases, the proposed method loses its advantage. In the Pyramids scene, all algorithms accurately reconstructed the convex and concave surfaces. The proposed method demonstrated some advantages in Stripe scene with lower demands for texture contrast. Table I presents the quantitative analyses of the scenes in Fig. 4 based on the BadPix(BP) and MSE metrics, along with the size of the network model parameters. Across various scenarios, the proposed method exhibited competitive advantages based on average quantitative metrics. Regarding the Average BP01, BP03, BP07 metrics, the proposed method shows lower misjudgment rates. Regarding the Network parameters metric, the proposed method is particularly efficient in lightweight design, having only 0.5M parameters, which makes it more suitable for resource-constrained environments than methods such as LLFNet and FastLFnet.
Although the proposed method maintains relatively low levels of BP and MSE metrics in specific scenes, such as Stripes, there is still room for further optimization, especially in handling complex textures or special structures. The proposed method demonstrates stable performance across different scenes, making it applicable to various scenarios.
@@ IV-B Algorithm analysis and evaluation using real data experiments
Real light field data typically contain more noise, non-Lambertian radiance data, and more complex occlusion scenarios. The Stanford Lytro light field dataset [46] was captured using a Lytro Illum camera. In this study, the proposed method was compared with the built-in algorithms of the Lytro camera, MV, LF_OCC, SPO, and AMN, as shown in Fig. 5.
@@
Figure 5: Disparity estimation results for data acquired using the Lytro Illum, with zoomed-in regions shown in close-up mode.
@@
In Fig. 5, the first two rows depict scenes with intricate occlusion structures suitable for evaluating the performance of the algorithm in handling complex scenarios. The proposed and other methods accurately predicted the intricate branches in the foreground. LF_OCC exhibited significant errors in the background region, SPO showed multiple protrusion errors at the boundaries of the branches, and the AMN and MV algorithms were smoothed out in detail, which can be observed in the magnified area indicated by the red box in the second row. In Fig. 5, the foreground in the last two rows is mainly comprised of iron fences. Apart from the built-in algorithm of the Lytro camera, both the proposed and other algorithms accurately reconstructed the iron fences in the foreground. However, in the background region, the proposed algorithm predicted the least noise, which was particularly notable in the scene reconstructed in the last row.
The experimental results demonstrate that the proposed method accurately predicts the disparity structure in real-world scenes and reconstructs more details in edge-occlusion areas.
IV-C Spatial super-resolution experiments
The proposed method parameterizes the continuous disparity, which provides a unique advantage because a discrete system matrix does not limit it. This allows the extraction of high-resolution disparity images from a well-trained network, thereby satisfying the need for super-resolution disparity image reconstruction. Figure 6 illustrates the differently sized reconstructed images extracted from the trained network using the Dino scene. This indicates that the quality of the disparity images was not compromised when extracting high-resolution images from the network.
Figure 6: Experimental results on super-resolution: Using the Dino from the 4D Light Field Benchmark [45]. (a) presents the reference view of the scene, and (b) displays the ground truth. (c) exhibits disparity reconstructed at resolutions of $512\times 512$, with (d) showing its profile curves at the 300th line. Panels (e) and (f) exhibit the disparity reconstructed at resolutions of $1024\times 1024$ and $2048\times 2048$, respectively.
IV-D Ablation studies
This section validates the design and parameter selection through an ablation study of the combination of loss functions and viewpoint selection loss strategy.
The ablation study results presented in Fig. 7 and Table II provides evidence for the rationale behind the proposed method of combining loss functions and designing loss functions for viewpoint selection. Table II presents the quantitative analysis results for the scenes depicted in Fig. 7. These scenes included a grid-cloth background, a pear with a bee pattern, and a wooden sphere with a black pattern. Analysis of the sphere’s disparity in the images reveals that using only the L1 loss leads to noticeable layering and rough surfaces on continuous surfaces. However, the predictions using both L1 and MSSIM losses resulted in smoother surfaces. Furthermore, the inclusion of TV regularization, which imposes sparse constraints on the gradient domain of the disparity, leads to smoother surfaces.
Figure 7: Ablation experiments using the Dino from HCI [44]. The first row is the disparity, whereas the second row (excluding (a)) is the profile curves. The ground truth curves are marked in red in the profile plots, and those reconstructed by the proposed method are shown in blue. In column (a), the first row is the ground truth and the second row is the reference view of the scene. (b) The results using the L1 norm; (c) the results using both L1 and MSSIM losses. (d) and (e) show the results using L1 and MSSIM losses along with TV regularization. Panels (b)-(d) use only half of the viewpoints for backpropagation, whereas panel (e) utilizes all viewpoints.
TABLE II: Quantitative analysis of ablation experiment, presented with BadPix0.03, BadPix0.05, BadPix0.07 and MSE ($\times$ 100) metrics (smaller scores are better).
StillLife MSE BadPix0.07 BadPix0.05 BadPix0.03
L1 0.87 10.42 21.07 43.92
L1+MSSIM 0.56 8.58 14.87 29.93
TV 0.52 5.01 8.46 21.50
Full view 1.55 7.28 9.60 16.32
As shown in Table II, the combination of MSSIM loss and TV constraint yielded the lowest MSE and BadPixel values, indicating a higher reconstruction accuracy. Figure 7 (d) applies a loss function based on viewpoint selection, transmitting only half of the viewpoints during each backpropagation step, whereas Fig. 7 (e) utilizes all the viewpoints. The profile curves in Fig. 7 demonstrate that, when using the complete set of viewpoints for backpropagation, the occlusion caused by the wings of the bee cannot be effectively removed. According to Table II, employing all viewpoints for backpropagation (full view row) results in higher values for MSE, BadPix0.07, and BadPix0.05 metrics compared to propagating only half of the viewpoints (TV row), indicating poorer reconstruction quality. In disparity reconstruction from light fields, occlusions and noise within a scene are the main sources of error, although these occluded and noisy regions account for a relatively small fraction of the overall scene. During the backpropagation process, the use of more viewpoints when propagating unoccluded areas improves the reconstruction accuracy. Consequently, regarding the BadPix0.03 metric, employing all available viewpoints for backpropagation (full view row) yielded lower values than using only half of the viewpoints (TV row), indicating superior reconstruction quality.
@@ V Discussion
We conducted comparative experiments on widely used light-field datasets to validate the performance of the proposed method in disparity estimation. The results demonstrate the exceptional proficiency of the proposed method in addressing inclined surfaces, smooth textureless regions, and occlusions in intricate scenes. When confronted with the weakly textured tail area of the butterfly and occluded leaf boundaries in the Papillon scene, the proposed method delivered a higher accuracy and smoother disparity estimates than MV, LF_OCC, SPO, and AMN. Further validation of the method’s resilience and adaptability was sought through tailored challenges in scenes such as backgammons, dots, pyramids, and stripes, which tested its ability to handle fine structures, noise interference, convex-concave geometries, and low-texture contrast scenarios. Despite a slight decline in the performance under high-noise conditions, the proposed method achieved effective disparity reconstruction and detail preservation in most cases. Additionally, experiments on the Stanford Lytro light field dataset demonstrated the success of the proposed method in accurately discerning disparity structures, even in highly noisy and complex occluded real-world scenes, particularly in suppressing background noise and reconstructing foreground edge details relative to contrastive algorithms. Moreover, by extracting disparities at varying resolutions directly from a pre-trained NDF, the proposed approach transcends the conventional limitations imposed by the reliance on discrete matrix-based disparity representations. Finally, the ablation studies shed light on the influence of combination strategies and selected viewpoints on the loss function. The MSSIM loss curbed the spread of local pixel errors. In contrast, including total variation regularization enforces constraints on the disparity gradient domain, ensuring smooth and coherent estimation outcomes. The capability to propagate with only half of the viewpoints bolsters the resistance to occlusion and noise.
Although the proposed method demonstrated excellent performance in the above experiments, the disparity map exhibited higher frequencies than the natural images, particularly in edge regions with disparity differences, resulting in oscillations in the predicted disparities. Thus, further research is required to investigate parameterized network models for disparities to enhance algorithm performance and applicability.
In summary, the proposed compact network model demonstrates efficiency and robustness in handling weak textures, varying noise levels, and occlusion scenarios. This could potentially provide more precise and reliable solutions for light field depth estimation in practical applications.
@@ VI Conclusions
We proposed a novel iterative approach to reconstruct the NDF that utilized neural network parameterization to achieve an implicit, continuous disparity function. The essential feature of this method is the implicit parameterization of the disparity by neural fields, which overcomes the resolution limitations inherent in traditional light-field disparity estimation. Compared to conventional methods based on discrete disparity maps, the proposed approach significantly reduced computational complexity while maintaining high precision and demonstrating superior robustness, particularly in handling complex scenes where occlusion and noise affect accurate disparity reconstruction. The proposed method employed a differentiable framework to construct a neural network model capable of modeling disparity continuously, thereby effectively avoiding the loss of detail caused by discrete representations. During the backpropagation optimization process, only a subset of rays was selected for propagation computation in the viewpoint direction, helping to mitigate occlusion effects and noise interference in the disparity estimation. Experimental validation indicates that the reconstructed NDF achieved accurate disparity on both synthetic and real datasets, enabling the effective recovery of high-resolution disparity.

### v2

@@ IV-A2 Specific scene experiments
To gain a comprehensive understanding of the proposed method’s performance, we set up isolated challenges in specific scenarios: backgammons, dots, pyramids, and stripes. The Backgammon scene evaluated the algorithm’s capability to handle fine structures and occlusion boundaries, while the Dots scene investigated the impact of varying levels of camera noise on the reconstruction of objects of different sizes. The Pyramids scene tested the algorithm’s performance on convex, concave, circular, and planar geometrical structures, and the Stripes scene specifically assessed the influence of texture and contrast on occlusion boundary recognition. We conduct comparative experiments with state-of-the-art methods, where LF_OCC, SPO, and AMN are optimization-based methods, OACC and SubFocal are supervised methods, and OPAL[34], PnPusp[31], and UnLFDisp[33] are unsupervised methods. The results are presented in Fig. 6.
@@
Figure 6: Visual comparison of disparity estimation methods across different scenes. From top to bottom, rows correspond to Backgammon, Dots, Pyramids, and Stripes scene. For each scene, the top row shows the estimated disparity, and the bottom row shows the corresponding MSE. Results (b)-(j) are derived from original data in [53].
@@
TABLE I: Quantitative comparisons of results on synthetic datasets, presented with BP01(BadPix0.01), BP03(BadPix0.03), BP07(BadPix0.07), MSE ($\times$ 100), and Q25 metrics.
Methods Optimization-based Supervised Unsupervised
LF_OCC SPO AMN Ours OACC SubFocal PnPusp OPAL UnLFDisp
Backgammon BP01 88.839 49.940 43.599 54.878 21.613 12.473 53.840 49.773 55.137
BP03 39.826 8.639 13.354 16.378 6.640 4.281 27.753 17.268 12.188
BP07 19.006 3.782 6.913 6.656 3.931 3.194 14.200 8.832 5.203
MSE 21.587 4.587 12.287 8.929 3.938 3.667 9.399 5.177 6.857
Q25 1.680 0.572 0.367 0.539 0.190 0.090 0.440 0.448 0.567
Dots BP01 65.762 58.079 21.658 4.538 21.022 15.511 72.861 87.631 54.987
BP03 23.552 35.068 18.428 4.485 3.040 1.524 48.081 68.839 3.383
BP07 5.822 16.274 18.428 4.380 1.510 0.899 29.864 48.309 1.531
MSE 3.301 5.238 2.936 5.433 1.418 1.301 4.769 7.442 1.435
Q25 0.577 0.973 0.003 0.003 0.198 0.134 0.905 2.100 0.762
Pyramids BP01 74.144 79.206 32.485 32.675 3.852 1.867 24.779 25.863 27.193
BP03 20.495 6.263 5.643 2.853 0.536 0.411 4.157 2.616 3.425
BP07 3.172 0.861 0.449 0.129 0.157 0.220 0.659 0.430 0.269
MSE 0.098 0.043 0.023 0.016 0.004 0.005 0.021 0.016 0.016
Q25 0.905 1.206 0.184 0.552 0.080 0.062 0.234 0.274 0.385
Stripes BP01 53.669 21.879 13.499 8.389 15.244 9.386 67.906 54.561 22.341
BP03 21.257 15.460 7.966 8.053 4.644 3.568 43.372 25.512 8.513
BP07 18.408 14.987 5.835 7.518 2.920 2.464 24.708 13.232 5.123
MSE 8.131 6.955 2.128 2.518 0.845 0.821 2.371 1.362 1.798
Q25 0.500 0.064 0.700 0.038 0.134 0.091 0.706 0.487 0.209
Average BP01 70.604 52.276 27.810 25.12 15.433 9.809 54.847 54.457 39.9145
BP03 26.283 16.358 11.348 7.942 3.715 2.446 29.591 28.559 6.877
BP07 11.602 8.976 7.906 4.671 2.130 1.694 17.358 17.701 3.032
MSE 8.279 4.206 4.344 4.224 1.551 1.449 4.140 3.499 2.527
Q25 0.916 0.704 0.313 0.283 0.151 0.094 0.571 0.827 0.481
Parameters (M) 0.159 5.01 5.06 5.12 1.047 59.1
@@
Figures 6 (b)-(j) correspond to the different algorithms. From the experimental results, it is evident that the proposed method successfully reconstructs disparities in various challenging scenarios. In the Backgammon scene, the proposed method performed remarkably well in handling complex textures, tilted surfaces, and textureless smooth regions, particularly when compared with LF_OCC, and SPO, although it failed to estimate disparities at the upper-left crevice accurately. In the Dots scene, the proposed method reconstructs more foreground points in the upper-left corner compared to the lower-right corner, indicating a certain robustness to noise; however, as the noise increases, the proposed method loses its advantage. In Pyramids scene, all algorithms accurately reconstructed the convex and concave surfaces. The proposed method demonstrated some advantages in Stripes scene with lower demands for texture contrast. Table I presents the quantitative analyses of the scenes depicted in Fig. 6 based on BadPix (BP) and mean squared error (MSE) metrics, along with the sizes of the network model parameters. Scores that exceed those of our method are indicated in bold. The proposed method exhibits certain advantages over optimization-based methods and demonstrates competitiveness against unsupervised methods, although it does not perform as well as supervised methods. Regarding the network parameters metric, the proposed method is particularly efficient in lightweight design, with only 0.159M trainable parameters and lower memory usage, making it more suitable for resource-constrained environments than methods such as OACC, SubFocal, and UnLFDisp.
@@ IV-B Algorithm analysis and evaluation using real data experiments
Real LF data typically contain more noise, non-Lambertian radiance, and complex occlusion scenarios. We evaluate the performance of our method on real LFs captured by a moving camera [51] and the Stanford Lytro LF dataset [50]. In this study, the proposed method is compared with POBR, LF_OCC, SPO, AMN, OACC, DistgDisp, SubFocal, and PnPusp, as shown in Fig. 7 and Fig. 8.
@@
Figure 7: Disparity estimation results for data from the Lytro Illum.
Figure 8: Disparity estimation results for data from the moving camera. The first row shows a Lego Bulldozer, and the third row shows complex geometry: Tarot Cards and Crystal Ball (large angular extent).
@@
In Fig. 7, the first two rows show scenes with complex occlusion structures, suitable for evaluating the algorithm’s performance in handling challenging scenarios. Both the proposed method and other methods accurately predict the intricate branches in the foreground. POBR exhibits over-smoothing, LF_OCC shows significant errors in the background, SPO introduces multiple protrusion errors at the branch boundaries, and AMN loses detail due to excessive smoothing. The supervised learning methods perform the best. In the last two rows of Fig. 7, the foreground mainly consists of iron fences, which are accurately reconstructed by both the proposed method and other algorithms. However, in the background, the proposed method produces the least noise compared to POBR, LF_OCC, SPO, AMN, and PnPusp, especially in the scene in the last row. In the third row of Fig. 7, the results from SubFocal show incorrect black spots within the gaps of the iron fence.
Fig. 8 presents the disparity estimation results for LF data captured using a moving camera. In the Lego Bulldozer scene, which contains very complex geometry and textureless areas, most algorithms struggle significantly with the textureless regions. The proposed method, unlike POBR, estimates more disparity in these textureless areas. In the Tarot Cards and Crystal Ball scene, the cards serve as diffuse-textured objects at various depths and angles, and the scene includes large disparities with a textureless Crystal Ball. In this challenging setup, PnPusp, OACC, and SubFocal perform the worst. While none of the algorithms accurately estimate the disparity of the Crystal Ball, the proposed method achieves the best performance.
The experimental results demonstrate that the proposed method effectively predicts disparity structures in real scenes for LF data captured by both Lytro and moving cameras, reconstructing more details compared to other methods.
IV-C Ablation studies
This section validates the effectiveness of the loss strategy in (6), where only half of the viewpoints are selected for backpropagation in each iteration. The effect of the TV constraint is also demonstrated. Fig. 9 presents the results of the ablation study. In Fig. 9, columns (b) and (c) share the same parameters except for the number of viewpoints used. Specifically, column (b) uses all viewpoints, leading to significant errors at the boundaries. In contrast, column (c) uses only half of the viewpoints as specified by (6), significantly reducing boundary errors. This improvement is reflected in both the MSE maps and the performance metrics. Comparing columns (c) and (d), column (c) applies the TV constraint, while column (d) does not. The reconstruction results show that, in non-boundary texture regions of the scene, the disparity surfaces in column (c) are smoother, whereas column (d) exhibits more noise in these areas, as evident in the error maps.
Figure 9: The first row shows the disparity maps, while the second row (excluding (a)) presents the MSE maps. Column (b) uses all views, while column (c) uses half of the views as specified in (6). Column (d) uses half of the views and does not use TV
@@ V Discussion
We conduct comparative experiments on multiple LF datasets to validate the performance of the proposed method in disparity estimation. The results demonstrate the effectiveness of the proposed method in addressing inclined surfaces, smooth textureless regions, and occlusions in intricate scenes. Specifically, when challenged with the weakly textured tail area of the butterfly and the occluded leaf boundaries in the Papillon scene, the proposed method achieves higher accuracy and smoother disparity estimates than BOPR, LF_OCC, SPO, and AMN. Further validation is obtained through targeted experiments in scenes such as backgammon, dots, pyramids, and stripes, designed to test its ability to handle fine structures, noise, convex-concave geometries, and low-texture contrast scenarios. Although a slight decline in performance is observed under high-noise conditions, the proposed method achieves effective disparity reconstruction and preserves detail in most cases.
Additionally, experiments on the Stanford Lytro and moving camera dataset demonstrate the proposed method’s ability to discern disparity structures in noisy and complex occluded real-world scenes, excelling in background noise suppression and foreground edge reconstruction compared to other algorithms. While there remains a disparity gap compared to supervised or unsupervised methods relying on large datasets, the proposed method reconstructs greater disparity detail in challenging scenes.
Disparity describes the geometric structure of a scene, and disparity maps typically exhibit higher frequencies than natural images, especially along edges with disparity discontinuities. As a result, the proposed method exhibits oscillations in these regions. Future research should explore parameterized network models tailored for disparity estimation to improve performance. In summary, the proposed compact network model shows efficiency and robustness in handling weak textures, varying noise levels, and occlusion scenarios, offering a solution for precise LF depth estimation in practical applications.
@@ VI Conclusions
We propose an iterative approach to reconstruct the NDF that utilizes neural network parameterization to achieve an implicit, continuous disparity function. The essential feature of this method is the implicit parameterization of the disparity by neural fields, which overcomes the resolution limitations inherent in traditional LF disparity estimation. Compared to conventional methods based on discrete disparity maps, the proposed approach reduces computational complexity while maintaining high precision and demonstrating superior robustness, particularly in handling complex scenes where occlusion and noise affect accurate disparity reconstruction. The proposed method employs a differentiable framework to construct a neural network model capable of modeling disparity continuously, thereby effectively avoiding the loss of detail caused by discrete representations. During the backpropagation optimization process, only a subset of rays is selected for propagation computation in the viewpoint direction, helping to mitigate occlusion effects and noise interference in the disparity estimation. Experimental validation demonstrates that the reconstructed NDF estimates disparity on both synthetic and real datasets.